140 g of S2Cl2 was collected from the reaction of 640 g of Sulfur and 142 g of Chlorine. What is the maximum amount of S2Cl2 in grams that can be produced from the reaction?


1
Expert's answer
2022-05-24T07:06:05-0400

Solution:

The molar mass of sulfur (S8) is 256 g/mol

The molar mass of chlorine gas (Cl2) is 71 g/mol


Calculate moles of each reactant:

Moles of S8 = (640 g S8) × (1 mol S8 / 256 g S8) = 2.5 mol S8

Moles of Cl2 = (142 g Cl2) × (1 mol Cl2 / 71 g Cl2) = 2.0 mol Cl2


Balanced chemical equation:

S8(l) + 4Cl2(g) → 4S2Cl2(l)

According to stoichiometry:

1 mol of S8 reacts with 4 mol of Cl2

Thus, 2.5 mol of S8 react with:

(2.5 mol S8) × (4 mol Cl2 / 1 mol S8) = 10 mol Cl2

However, initially there is 2.0 mol of Cl2 (according to the task)

Therefore, Cl2 acts as limiting reagent and S8 is excess reagent


According to stoichiometry:

4 mol of Cl2 produce 4 mol of S2Cl2

Thus, 2.0 mol of Cl2 produce:

(2.0 mol Cl2) × (4 mol S2Cl2 / 4 mol Cl2) = 2.0 mol S2Cl2


The molar mass of S2Cl2 is 135 g/mol

Therefore,

Mass of S2Cl2 = (2.0 mol S2Cl2) × (135 g S2Cl/ 1 mol S2Cl2) = 270 g S2Cl2

Mass of S2Cl2 = 270 g


Answer: The maximum amount of S2Clthat can be produced from the reaction is 270 grams